OpenStudy (anonymous):

The point (1,4) lies on the graph of an equation y=f(x) for which dy/dx=4 sqrt(xy) where x>= 0 and y >= 0. When x = 0 the value of y is:

OpenStudy (anonymous):

I was thinking of doing a differential equation setup, is this the right approach?

OpenStudy (anonymous):

Something like this: \[\int\limits_{}^{}dy \div \sqrt{y} = \int\limits_{}^{}\sqrt{x}dx\]

OpenStudy (freckles):

you are missing 4 but close

OpenStudy (freckles):

\[\int\limits\limits_{}^{}dy \div \sqrt{y} = \int\limits\limits_{}^{}4 \sqrt{x}dx\]

OpenStudy (freckles):

Have you integrated yet?

OpenStudy (freckles):

you are suppose to find y by evaluating the above then find the constant by using the point (!,4) then you can find y for whatever x they want where x>0 of course

OpenStudy (anonymous):

Thanks @freckles !! How could I forget something as silly as that 4...and to think I had even put the 4 onto my paper.. I completely forgot to include it in the integration. Thank you very much

OpenStudy (freckles):

what did you get for C?

OpenStudy (anonymous):

I got C = 2/3 but I'm starting to see this answer is wrong..

OpenStudy (freckles):

yep it is a bit off